OpenStudy (anonymous):

Which value of x satisfies both -9x + 4y = 8 and -3x − y = 4?

OpenStudy (anonymous):

@Kemist helppppppppppppp

OpenStudy (anonymous):

Choices?

OpenStudy (anonymous):

 @Kemist

OpenStudy (anonymous):

E? @Kemist

OpenStudy (anonymous):

Plug in the numbers to (x) until both equations are equal. :)

OpenStudy (anonymous):

I give up im just guessing screw the test.

OpenStudy (anonymous):

No, I'll help you

OpenStudy (anonymous):

-9(-1/9)+4y=8 -3(-1/9)-y=4 Solve this

OpenStudy (anonymous):

no i freaking have 20 minutes now. and i still have 49 more questions and i dont know ANY of them. # screw summerschool #

OpenStudy (anonymous):

@Kemist i DO NOT know how to do any of this ok!??

OpenStudy (anonymous):

Lol I don't understand it half the time either. But here @efco_chen he's on the same kind of problems. I'm sure he can help you more than I can.

OpenStudy (anonymous):

im just gunna start putting a on everything i legit give up

OpenStudy (anonymous):

My teacher says C

OpenStudy (anonymous):

really?